OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = x − 2, find f[g(−6)].

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

f(g(x))=(g(x)-4)=((x-2)-4)

OpenStudy (anonymous):

f[g(−6)]=[g(6)]-4=(x-2)-4

OpenStudy (anonymous):

Solve for f(g(x)) and then put in x.

OpenStudy (anonymous):

it would be 4-(-6) then g=(-6)-2

OpenStudy (anonymous):

then you sove that? @Numb3r1

OpenStudy (anonymous):

if you got the value of g(-6) make it an x value and plug it in the function of f(x) making it f(g(-6))

OpenStudy (anonymous):

Yeah, you solve (x-2)-4 with x=-6.